# Machine-precision energy conservative quadrature hyperreduction of Lagrangian hydrodynamics

### v1

@@ A.2 Conservative EQP
In this section we offer details on the implementation of the NNLS problem for the energy conservative EQP method developed in Section 4 and on an additional basis enrichment operation performed to ensure conservation of the discrete total energy for simulations consisting of multiple time windows.
@@
Following the approach outlined in Section 3.1 and A.1 the reduced quadrature rule is computed as the solution of an NNLS problem with a total of $N_{c}=(n_{v}+n_{e})N_{t}$ accuracy constraints, where $N_{t}$ denotes the number of snapshots of $\tilde{w}$ used to form the constraints. The constraints are used to control the accuracy of a single reduced quadrature rule that will be used for the approximation of both force vectors $\bm{F}_{v}^{\phi}$ and $\bm{F}_{e}^{\phi}$ introduced in (3.10)
$\bm{F}_{v}^{\phi}(\tilde{\bm{w}})=\int_{\Omega(t)}\bm{g}_{v}^{\phi}(x,\tilde{w})dx\qquad\bm{F}_{e}^{\phi}(\tilde{\bm{w}},\tilde{\bm{v}}^{\prime})=\int_{\Omega(t)}\bm{g}_{e}^{\phi}(x,\tilde{w},\tilde{v})dx$
@@
$g_{v,i}^{\phi}(x,\tilde{w})=\sigma(\tilde{w}):\nabla\phi^{v}_{i}dx\qquad g_{e,i}^{\phi}(x,\tilde{w},\tilde{v})=(\sigma(\tilde{w}):\nabla\tilde{v})\phi^{e}_{i}dx.$
The accuracy constraints are encoded in a matrix $C\in\mathbb{R}^{N_{c}\times J}$ where the first $n_{v}N_{t}$ rows are used to enforce the constraints related to $\bm{F}_{v}^{\phi}$
@@
with constraint index $s=n_{v}N_{t}+i+kn_{e}$. In the above $\tilde{w}(t_{k})$ denotes the snapshots used to form the constraints.
@@
When the simulation consists of more than one time window, an additional modification is made to the reduced bases $\bm{\Phi}_{v}$ and $\bm{\Phi}_{e}$ to ensure that the total energy is conserved across windows. Let $\tilde{\bm{v}}^{(n-1)}$ and $\tilde{\bm{e}}^{(n-1)}$ denote the velocity and energy solution vectors after the final timestep of window n1n-1, $n>1$ has been completed
@@
Note that no timestepping has been performed after switching from window n1n-1 to $n$, so the two pairs of vectors should be identical. To ensure that this is the case we respectively extend bases $\bm{\Phi}_{v}^{(n)}$ and $\bm{\Phi}_{e}^{(n)}$ by adding the solution vectors $\tilde{\bm{v}}^{(n-1)}$ and $\tilde{\bm{e}}^{(n-1)}$ as an additional column and reorthonormalize. This guarantees that $\tilde{\bm{v}}^{(n-1)}$ and $\tilde{\bm{e}}^{(n-1)}$ are in the span of the respective bases $\bm{\Phi}_{v}^{(n)}$ and $\bm{\Phi}_{e}^{(n)}$. In turn, this ensures that energy is conserved exactly when changing windows
@@
No such extension is required for the first window ($n=0$). Since the extension of the bases $\bm{\Phi}_{v}^{(n)}$ and $\bm{\Phi}_{e}^{(n)}$ requires knowledge of the state vectors $\tilde{\bm{v}}^{(n-1)}$ and $\tilde{\bm{e}}^{(n-1)}$, this operation must be performed during the online stage at the time of the window change. To extend and incrementally reorthonormalize the bases efficiently we employ the modified Gram-Schmidt algorithm with double orthogonalization [88].

### v2

@@ A.2 Conservative EQP
We consider the implementation of the NNLS problem for the energy conservative EQP method developed in Section 4, as well as an additional basis enrichment operation performed to ensure conservation of the discrete total energy for simulations consisting of multiple time windows.
@@
Following the approach outlined in Section 3.1 and A.1, the reduced quadrature rule is computed by solving an NNLS problem with a total of $N_{c}=(n_{v}+n_{e})N_{t}$ accuracy constraints, where $N_{t}$ denotes the number of snapshots of $\tilde{w}$ used to form the constraints. The constraints are used to control the accuracy of a single reduced quadrature rule that will be used for the approximation of both force vectors $\bm{F}_{v}^{\phi}$ and $\bm{F}_{e}^{\phi}$ introduced in (3.10)
$\bm{F}_{v}^{\phi}(\tilde{\bm{w}})=\int_{\Omega(t)}\bm{g}_{v}^{\phi}(x,\tilde{w})dx\qquad\bm{F}_{e}^{\phi}(\tilde{\bm{w}},\tilde{\bm{v}})=\int_{\Omega(t)}\bm{g}_{e}^{\phi}(x,\tilde{w},\tilde{v})dx$
@@
$g_{v,i}^{\phi}(x,\tilde{w})=\sigma(\tilde{w})(x):\nabla\phi^{v}_{i}(x)\qquad g_{e,i}^{\phi}(x,\tilde{w},\tilde{v})=(\sigma(\tilde{w})(x):\nabla\tilde{v}(x))\,\phi^{e}_{i}(x).$
The accuracy constraints are encoded in the matrix $\bm{C}\in\mathbb{R}^{N_{c}\times J}$, where the first $n_{v}N_{t}$ rows are used to enforce the constraints related to $\bm{F}_{v}^{\phi}$
@@
with constraint index $s=n_{v}N_{t}+i+kn_{e}$. In the above, $\tilde{w}(t_{k})$ denotes the snapshots used to form the constraints.
@@
When the simulation consists of more than one time window, an additional modification is made to the reduced bases $\bm{\Phi}_{v}$ and $\bm{\Phi}_{e}$ to ensure that the total energy is conserved across windows. Let $\tilde{\bm{v}}^{(n-1)}$ and $\tilde{\bm{e}}^{(n-1)}$ denote the velocity and energy solution vectors after the final timestep of window n1n-1, $n>1$, has been completed
@@
Note that no timestepping has been performed after switching from window n1n-1 to $n$, so the two pairs of vectors should be identical. To ensure that this is the case, we extend the basis matrices $\bm{\Phi}_{v}^{(n)}$ and $\bm{\Phi}_{e}^{(n)}$ by adding the respective solution vectors $\tilde{\bm{v}}^{(n-1)}$ and $\tilde{\bm{e}}^{(n-1)}$ as an additional column and reorthonormalize. This guarantees that $\tilde{\bm{v}}^{(n-1)}$ and $\tilde{\bm{e}}^{(n-1)}$ are in the span of the respective basis matrices $\bm{\Phi}_{v}^{(n)}$ and $\bm{\Phi}_{e}^{(n)}$. In turn, this ensures that energy is conserved exactly when changing windows
@@
No such extension is required for the first window ($n=0$). Since the extension of the basis matrices $\bm{\Phi}_{v}^{(n)}$ and $\bm{\Phi}_{e}^{(n)}$ requires knowledge of the state vectors $\tilde{\bm{v}}^{(n-1)}$ and $\tilde{\bm{e}}^{(n-1)}$, this operation must be performed during the online stage at the time of the window change. To extend and incrementally reorthonormalize the basis matrices efficiently, we employ the modified Gram-Schmidt algorithm with double orthogonalization [71].